# Accelerating Edge Inference for Distributed MoE Models with Latency-Optimized Expert Placement

### v2

@@ Abstract
Mixture-of-Experts (MoE) have become a cornerstone for training and scaling large language models (LLMs), offering substantial gains in model capacity and efficiency through sparse expert activation. However, serving these models remains challenging in practice, particularly in resource-constrained edge environments, due to their large memory footprint and complex communication demands. While centralized cloud inference is common, it incurs high infrastructure costs, along with latency and privacy concerns. A few recent edge MoE works propose memory-efficient strategies but typically focus on single-device or homogeneous setups. This paper presents DanceMoE, an efficient MoE inference framework that enables activation-aware expert placement across collaborative, heterogeneous, GPU-equipped edge servers. DanceMoE leverages the inherent sparsity of MoE models and workload locality to minimize cross-server communication and enable efficient expert placement under heterogeneous resource constraints. It introduces a data-driven, activation-aware placement algorithm that balances local coverage and memory usage across servers, alongside a lightweight migration mechanism that adapts expert assignments under evolving workloads. We evaluate DanceMoE on modern MoE models and widely used datasets, demonstrating up to 30.6% lower inference latency, and substantial communication reduction compared to state-of-the-art baselines, showcasing the effectiveness of collaborative edge-based MoE inference.
@@ I Introduction
Mixture-of-Experts (MoE) architectures have emerged as a central design choice for training large language models (LLMs) at scale. By integrating multiple specialized subnetworks, known as experts, MoE achieves substantial performance gains without a proportional increase in computational cost. A lightweight gating mechanism dynamically routes each input token to only a small subset of experts, enabling sparse activation and efficient parallelism during training. This sparsity enables MoE to scale significantly while keeping training costs manageable, leading to its adoption in state-of-the-art models such as Switch Transformer [fedus2022switch], Mixtral [jiang2024mixtral], and DeepSeek-V3 [liu2024deepseek]. As a result, MoE has become a foundational architecture for training and scaling modern LLMs.
@@
Despite its efficiency during training, MoE inference remains resource-intensive, especially in terms of GPU memory, due to growing sizes of modern MoE models. Even with sparse activation, the cumulative memory footprint often exceeds the capacity of a single GPU. For example, Mixtral-8×7B [jiang2024mixtral] requires over 80 GB of GPU memory, far exceeding the 16–24 GB typically available on edge servers with commodity GPUs, such as those equipped with NVDIA RTX 4090 or A4000 GPUs commonly found in university labs, hospitals, small and medium-sized enterprises [nvdia4090, gpuion]. As a result, the standard practice is to offload inference to cloud servers with large-scale GPU clusters. However, relying solely on centralized cloud services presents many limitations: prohibitively high expenditures of renting cloud GPU instances, high response latency from network round-trips, and growing concerns over data privacy and regulatory compliance.
@@
Why collaborative edge inference for MoE? In response, distributed deployments of LLMs including MoE models on GPU-equipped edge servers has emerged as a compelling alternative. Many organizations (e.g., universities and research labs) coordinate underutilized lab servers or desktops with modest GPU capabilities in different locations [gpuion]. By doing so, edge-based inference offers significantly lower infrastructure costs and improved data locality. However, these methods still incur high MoE inference latency, as they lack fine-grained analysis of MoE model architectures and expert activation patterns. A few recent works have proposed strategies for edge MoE, but they either target acceleration on a single edge device [edgemoe] or overlook the system and request heterogeneity in edge scenarios [fast-edge-moe, go2025moetuner]: resource configurations, request volumes, and expert usages can all vary across distributed edge servers. To address this, we propose a collaborative MoE inference architecture, inspired by Expert Parallelism (EP), which distributes each layer of experts across GPUs. As illustrated in Fig. 1, inference is performed collaboratively across multiple devices, with each server hosting only a subset of the experts due to memory constraints. When a request arrives at Server 1 that activates an expert not present locally, such as Expert 8, the computation can still be completed by invoking the expert remotely from another server like Server 2 or Server 3. This setup leverages idle edge resources to form a cooperative inference layer, mitigating memory bottlenecks without relying on cloud execution.
@@
Importance and challenges of expert placement. Despite its advantages, collaborative MoE edge inference still faces challenges, particularly inter-GPU/server communication latency. Existing distributed MoE inference methods [shoeybi2019megatron, go2025moetuner, deepseek2025eplb] including EP solutions designed for datacenters are unsuitable in edge scenarios due to two reasons: (i) edge servers vary in GPU memory, compute capacity, and network reliability, unlike datacenters with uniform hardware and high-speed interconnects; and (ii) redirecting requests across locations is costly, making it essential to process in proximity to data, a core principle of edge computing. These constraints call for strategic expert placement to reduce inference latency by maximizing local execution and minimizing remote calls. Additionally, expert activation patterns vary across servers due to user behavior and workload differences. While stable over short periods, longer-term shifts can degrade performance if placements remain fixed. Lightweight, periodical placement adjustment is needed to adapt layouts as workloads evolve, without complex migration protocols or frequent re-optimization.
@@
Insights. Fortunately, the inherent sparsity and modularity of MoE models offer an opportunity to address these challenges. Since only a small subset of experts is activated per input, systems can be designed to favor local execution. As explored in Section II-A, servers often exhibit distinct patterns of expert activation due to varying user tasks and input distributions, allowing each node to prioritize caching its most frequently accessed experts. As shown in Fig. 1, Server 1 and Server 2 each store only four out of eight experts, strategically selecting those most relevant to their local workloads. Consequently, Server 2 can process inputs routed to Experts 4 and 7 entirely locally, while Server 1 incurs occasional cross-server communication. Crucially, such remote calls remain infrequent, keeping end-to-end latency within acceptable bounds. Leveraging these insights, we propose DanceMoE, a novel Distributed activation-aware collaborative MoE inference system tailored to resource-constrained edge environments. Our approach explicitly addresses the challenges of memory limitations, communication overhead, and workload heterogeneity, while capitalizing on the sparse activation patterns intrinsic to MoE architectures. The key contributions of this work are:
@@
    System Design. We introduce DanceMoE, a novel distributed inference system for collaborative MoE serving across memory-constrained, heterogeneous edge servers. It exploits activation sparsity and local workload patterns to minimize cross-server communication and enable scalable deployment.
@@
    Lightweight Expert Migration. To adapt to evolving workloads and input distributions, DanceMoE periodically re-evaluates placement and triggers efficient migrations when beneficial—without incurring high coordination overhead.
@@
    Comprehensive Experiments. We evaluate DanceMoE using multiple MoE models and real-world data. Compared to existing approaches, DanceMoE reduces cross-server communication and achieves up to 30.6% lower inference latency.
@@ II-A Expert Activation Patterns
A core property of MoE models is their sparse activation: only a small subset of experts is activated per layer for each input. While expert activations may appear evenly distributed across general-purpose tasks, we make a key observation: for specific task types, the activation patterns become highly skewed and task-dependent. As illustrated in Fig. 2, when the Mixtral [jiang2024mixtral] model processes arithmetic tasks versus ASCII word recognition tasks from BIG-bench dataset [srivastava2023beyond], the expert utilization at the first layer varies significantly. Arithmetic tasks exhibit a strongly skewed activation pattern where Expert 0 dominates, while Expert 1 is rarely used. In contrast, ASCII word recognition predominantly activates Expert 3.
@@ II-B The Need of Collaborative Inference
As discussed above, one way to accommodate low-frequency experts is through memory offloading. For instance, MoE-Infinity [moe-infinity] enables single-device MoE inference by dynamically loading rarely used experts from RAM to GPU memory, while frequently activated experts remain cached on the GPU. However, such methods are limited to a single-machine setup and cannot take advantage of the collective computing power available in edge environments. In practice, edge servers often exhibit heterogeneous workloads and resource capacities. For example, Mooncake’s open-source online trace dataset [qin2025mooncake] shows that ToolAgent requests occur much more frequently than conversational queries. Moreover, the hardware heterogeneity among servers, such as differences in GPU availability or compute capability, leads to resource imbalance: some nodes remain underutilized while others are bottlenecked. To address this, a natural idea is to redistribute the load across servers.
One straightforward solution is to redirect requests to idle servers, i.e., reroute incoming requests to another server that performs inference locally using memory offloading. While this improves load balancing to some extent, our experiments show that it remains suboptimal. To highlight this, we compare three methods in Table I. The experiment uses the Mixtral-8×7B model deployed across three simulated edge servers. Each server processes a distinct type of request drawn from three datasets: arithmetic reasoning, ASCII word recognition, and abstract narrative understanding (from BIG-bench [srivastava2023beyond]). MoE-Infinity (w/ LB) represents the request-redirection baseline. In contrast, Naive Collaboration deploys experts randomly across the servers and enables distributed inference with remote expert calls. Despite its simplicity, the collaborative setup achieves significantly lower average latency, showing more balanced load across all servers.
@@ III System Design
Figure 4: The workflow of DanceMoE. The system consists of two primary components working in coordination to enable efficient distributed inference for MoE models: a global scheduler and a runtime multi-server system that executes inference.
@@ 22          $N_{l,\text{total}}\leftarrow N_{l,\text{total}}+1$;
This follows from the asymptotic equipartition property [Cover]: the typical set covering $(1-\delta)$ of the probability mass has cardinality lower bounded by $2^{H(\mathbf{p})-\delta\log E}$. ∎
@@ IV Implementation and Experimental Results
Based on the open-source code of MoE-infinity [moe-infinity], we implement DanceMoE. Building upon the MoE-Infinity framework, we have developed a multi-server MoE inference system. This flexible architecture allows us to accommodate various hardware configurations and optimize resource utilization to meet the demands of various deployment scenarios.
@@ IV-A Experimental Setup
MoE Model. Our experiments employ two MoE architectures: (1) Mixtral-8x7B [jiang2024mixtral] (32 layers with 2 active experts out of 8 per layer), whose complete parameter set exceeds single-GPU memory capacity, and (2) Deepseek-V2-Lite [liu2024deepseekv2] (26 layers selecting 8 active experts from 64 per layer, totaling 1,664 possible expert combinations), which presents greater routing complexity while being more memory-efficient. To simulate realistic edge deployment scenarios, we artificially constrain GPU memory allocation to 70% of total capacity for Mixtral-8x7B and 30% for Deepseek-V2-Lite, reflecting their respective memory requirements and operational constraints.
Dataset. Our experimental evaluation incorporated four widely adopted benchmark datasets: BIG-bench [srivastava2023beyond], MMLU-Pro [wang2024mmlu], WikiText [merity2016pointer], and Tako [li2023taco].
@@
  • Uniform. Experts are uniformly distributed across all available devices. For instance, in our experimental setup with 4 GPUs, each device hosts 2 random experts per layer for the Mixtral model and 16 experts per layer for Deepseek-V2-Lite. This placement method aligns with the expert parallelism implementation in the Megatron-LM [shoeybi2019megatron] project.
@@
  • SmartMoE [zhai2023smartmoe]. While SmartMoE’s core focus is training optimization, our framework specifically implements SmartMoE’s placement module, which strategically distributes experts across GPUs based on real-time workload. The placement algorithm balances workload distribution to maintain uniform expert allocation across devices.
  • Expert Parallelism Load Balancer (EPLB) [deepseek2025eplb]. EPLB is a redundant expert strategy in DeepSeek-V3 [liu2024deepseek] that duplicates high-load experts and heuristically distributes them to balance GPU workloads. Since the open-source implementation only supports homogeneous systems, we re-implement the algorithm within our framework to accommodate heterogeneous system configurations.
@@ IV-B System Performance
Inference Latency. As shown in Tab. II, robustness across datasets and server configurations highlights practical superiority of DanceMoE. Uniform and SmartMoE achieve limited success due to inefficient resource utilization, while Redundance, though introducing expert duplication, suffers from higher latency due to suboptimal implementation. On the BigBench dataset, DanceMoE reduces the average latency by 30.6% compared to EPLB, the second-best method for DeepSeek-V2-Lite. For Mixtral, it achieves a slight but consistent improvement over EPLB. On the MultiData dataset, DanceMoE demonstrates a 10.9% reduction in latency compared to EPLB for DeepSeek-V2-Lite and a 2.0% improvement for Mixtral. Since EPLB is a load-balancing algorithm that does not account for cross-machine communication reduction, it underperforms compared to DanceMoE. For the Mixtral model, where the number of experts is relatively small and server capacity is limited, the room for optimization is narrower, resulting in smaller performance gains for DanceMoE.
Local Compute Ratio. To assess the impact of cross-machine communication on inference latency, we measured the local computation ratio as an indicator of communication overhead. Except for the Uniform and Redundance baselines, all methods adopt DanceMoE’s migration strategy but differ in their expert placement strategies. As shown in Fig. 6, DanceMoE consistently achieves higher local computation ratios. Notably, the advantage becomes more pronounced following explicit migration events triggered at 5 minutes (BigBench) and 10 minutes (MultiData).
@@ IV-C Large-scale Simulation
Simulation Setup. The simulator’s primary function orchestrates the inference process by: selecting the next layer for processing; updating server timestamps; recording completion times; computing inference latency. Through incremental time accumulation, it accurately models the temporal progression of the entire system. The simulator contains several key components: (1) Prompt Sequence Generator: We integrate a Poisson arrival time sequence with operational data collected from DanceMoE, the dataset contains both expert selection patterns and token processing volumes. (2) Prompt Routing Generator: To simulate data flow under varying expert placement configurations. (3) Communication & Computation Time Consumption Estimator: We develop a linear model to predict processing time per token batch. (4) Time Stamp Calculator. This module captures both communication and computation events, timestamping each operation. (5) System Timeline Scheduler. The scheduler carefully interleaves communication and computation events according to our prescribed design.

### v3

@@ Abstract
The emergence of Mixture-of-Experts (MoE) has transformed the scaling of large language models by enabling vast model capacity through sparse activation. Yet, converting these performance gains into practical edge deployment remains difficult, as the massive memory footprint and communication demands often overwhelm resource-limited environments. While centralized cloud-based solutions are available, they are frequently plagued by prohibitive infrastructure costs, latency issues, and privacy concerns. Moreover, existing edge-oriented optimizations largely overlook the complexities of heterogeneous hardware, focusing instead on isolated or uniform device setups. In response, this paper proposes Prism, an inference framework engineered for collaborative MoE serving across diverse GPU-equipped edge servers. By leveraging the intrinsic sparsity and input locality of MoE workloads, Prism minimizes inter-server communication and optimizes expert placement within diverse resource constraints. The framework integrates an activation-aware placement strategy that balances local request coverage with memory utilization, supplemented by a runtime migration mechanism to adapt expert distribution to dynamic workload changes. Experiments on contemporary MoE models and datasets demonstrate that Prism reduces inference latency by up to 30.6% and significantly lowers communication costs compared to state-of-the-art baselines, confirming the effectiveness of cooperative edge-based MoE serving.
@@ I Introduction
Mixture-of-Experts (MoE) architectures have emerged as a central design choice for training large language models (LLMs) at scale. By integrating multiple specialized subnetworks, known as experts, MoE achieves substantial performance gains without a proportional increase in computational cost. A lightweight gating mechanism dynamically routes each input token to only a small subset of experts, enabling sparse activation and efficient parallelism during training. This sparsity enables MoE to scale significantly while keeping training costs manageable, leading to its adoption in state-of-the-art models such as Switch Transformer [6], Mixtral [11], and DeepSeek-V3 [21]. As a result, MoE has become a foundational architecture for training and scaling modern LLMs.
@@
Despite its efficiency during training, MoE inference remains resource-intensive, especially in terms of GPU memory, due to growing sizes of modern MoE models. Even with sparse activation, the cumulative memory footprint often exceeds the capacity of a single GPU. For example, Mixtral-8×7B [11] requires over 80 GB of GPU memory, far exceeding the 16–24 GB typically available on edge servers with commodity GPUs, such as those equipped with NVDIA RTX 4090 or A4000 GPUs commonly found in university labs, hospitals, small and medium-sized enterprises [24, 19]. As a result, the standard practice is to offload inference to cloud servers with large-scale GPU clusters. However, relying solely on centralized cloud services presents many limitations: prohibitively high expenditures of renting cloud GPU instances, high response latency from network round-trips, and growing concerns over data privacy and regulatory compliance.
@@
Why collaborative edge inference for MoE? In response, distributed deployments of LLMs including MoE models on GPU-equipped edge servers has emerged as a compelling alternative. Many organizations (e.g., universities and research labs) coordinate underutilized lab servers or desktops with modest GPU capabilities in different locations [19]. By doing so, edge-based inference offers significantly lower infrastructure costs and improved data locality. However, these methods still incur high MoE inference latency, as they lack fine-grained analysis of MoE model architectures and expert activation patterns. A few recent works have proposed strategies for edge MoE, but they either target acceleration on a single edge device [34] or overlook the system and request heterogeneity in edge scenarios [5, 7]: resource configurations, request volumes, and expert usages can all vary across distributed edge servers. To address this, we propose a collaborative MoE inference architecture, inspired by Expert Parallelism (EP), which distributes each layer of experts across GPUs. As illustrated in Fig. 1, inference is performed collaboratively across multiple devices, with each server hosting only a subset of the experts due to memory constraints. When a request arrives at Server 1 that activates an expert not present locally, such as Expert 8, the computation can still be completed by invoking the expert remotely from another server like Server 2 or Server 3. This setup leverages idle edge resources to form a cooperative inference layer, mitigating memory bottlenecks without relying on cloud execution.
@@
Importance and challenges of expert placement. Despite its advantages, collaborative MoE edge inference still faces challenges, particularly inter-GPU/server communication latency. Existing distributed MoE inference methods [28, 7, 3] including EP solutions designed for datacenters are unsuitable in edge scenarios due to two reasons: (i) edge servers vary in GPU memory, compute capacity, and network reliability, unlike datacenters with uniform hardware and high-speed interconnects; and (ii) redirecting requests across locations is costly, making it essential to process in proximity to data, a core principle of edge computing. These constraints call for strategic expert placement to reduce inference latency by maximizing local execution and minimizing remote calls. Additionally, expert activation patterns vary across servers due to user behavior and workload differences. While stable over short periods, longer-term shifts can degrade performance if placements remain fixed. Lightweight, periodical placement adjustment is needed to adapt layouts as workloads evolve, without complex migration protocols or frequent re-optimization.
@@
Insights. Fortunately, the inherent sparsity and modularity of MoE models offer an opportunity to address these challenges. Since only a small subset of experts is activated per input, systems can be designed to favor local execution. As explored in Section II-A, servers often exhibit distinct patterns of expert activation due to varying user tasks and input distributions, allowing each node to prioritize caching its most frequently accessed experts. As shown in Fig. 1, Server 1 and Server 2 each store only four out of eight experts, strategically selecting those most relevant to their local workloads. Consequently, Server 2 can process inputs routed to Experts 4 and 7 entirely locally, while Server 1 incurs occasional cross-server communication. Crucially, such remote calls remain infrequent, keeping end-to-end latency within acceptable bounds. Leveraging these insights, we propose Prism, a novel Distributed activation-aware collaborative MoE inference system tailored to resource-constrained edge environments. Our approach explicitly addresses the challenges of memory limitations, communication overhead, and workload heterogeneity, while capitalizing on the sparse activation patterns intrinsic to MoE architectures. The key contributions of this work are:
@@
    System Design. We introduce Prism, a novel distributed inference system for collaborative MoE serving across memory-constrained, heterogeneous edge servers. It exploits activation sparsity and local workload patterns to minimize cross-server communication and enable scalable deployment.
@@
    Lightweight Expert Migration. To adapt to evolving workloads and input distributions, Prism periodically re-evaluates placement and triggers efficient migrations when beneficial—without incurring high coordination overhead.
@@
    Comprehensive Experiments. We evaluate Prism using multiple MoE models and real-world data. Compared to existing approaches, Prism reduces cross-server communication and achieves up to 30.6% lower inference latency.
@@ II-A Expert Activation Patterns
A core property of MoE models is their sparse activation: only a small subset of experts is activated per layer for each input. While expert activations may appear evenly distributed across general-purpose tasks, we make a key observation: for specific task types, the activation patterns become highly skewed and task-dependent. As illustrated in Fig. 2, when the Mixtral [11] model processes arithmetic tasks versus ASCII word recognition tasks from BIG-bench dataset [1], the expert utilization at the first layer varies significantly. Arithmetic tasks exhibit a strongly skewed activation pattern where Expert 0 dominates, while Expert 1 is rarely used. In contrast, ASCII word recognition predominantly activates Expert 3.
@@ II-B The Need of Collaborative Inference
As discussed above, one way to accommodate low-frequency experts is through memory offloading. For instance, MoE-Infinity [33] enables single-device MoE inference by dynamically loading rarely used experts from RAM to GPU memory, while frequently activated experts remain cached on the GPU. However, such methods are limited to a single-machine setup and cannot take advantage of the collective computing power available in edge environments. In practice, edge servers often exhibit heterogeneous workloads and resource capacities. For example, Mooncake’s open-source online trace dataset [25] shows that ToolAgent requests occur much more frequently than conversational queries. Moreover, the hardware heterogeneity among servers, such as differences in GPU availability or compute capability, leads to resource imbalance: some nodes remain underutilized while others are bottlenecked. To address this, a natural idea is to redistribute the load across servers.
One straightforward solution is to redirect requests to idle servers, i.e., reroute incoming requests to another server that performs inference locally using memory offloading. While this improves load balancing to some extent, our experiments show that it remains suboptimal. To highlight this, we compare three methods in Table I. The experiment uses the Mixtral-8×7B model deployed across three simulated edge servers. Each server processes a distinct type of request drawn from three datasets: arithmetic reasoning, ASCII word recognition, and abstract narrative understanding (from BIG-bench [1]). MoE-Infinity (w/ LB) represents the request-redirection baseline. In contrast, Naive Collaboration deploys experts randomly across the servers and enables distributed inference with remote expert calls. Despite its simplicity, the collaborative setup achieves significantly lower average latency, showing more balanced load across all servers.
@@ III System Design
Figure 4: The workflow of Prism. The system consists of two primary components working in coordination to enable efficient distributed inference for MoE models: a global scheduler and a runtime multi-server system that executes inference.
@@ 22          $N_{l,\text{total}}\leftarrow N_{l,\text{total}}+1$;
This follows from the asymptotic equipartition property [2]: the typical set covering $(1-\delta)$ of the probability mass has cardinality lower bounded by $2^{H(\mathbf{p})-\delta\log E}$. ∎
@@ IV Implementation and Experimental Results
Based on the open-source code of MoE-infinity [33], we implement Prism. Building upon the MoE-Infinity framework, we have developed a multi-server MoE inference system. This flexible architecture allows us to accommodate various hardware configurations and optimize resource utilization to meet the demands of various deployment scenarios.
@@ IV-A Experimental Setup
MoE Model. Our experiments employ two MoE architectures: (1) Mixtral-8x7B [11] (32 layers with 2 active experts out of 8 per layer), whose complete parameter set exceeds single-GPU memory capacity, and (2) Deepseek-V2-Lite [20] (26 layers selecting 8 active experts from 64 per layer, totaling 1,664 possible expert combinations), which presents greater routing complexity while being more memory-efficient. To simulate realistic edge deployment scenarios, we artificially constrain GPU memory allocation to 70% of total capacity for Mixtral-8x7B and 30% for Deepseek-V2-Lite, reflecting their respective memory requirements and operational constraints.
Dataset. Our experimental evaluation incorporated four widely adopted benchmark datasets: BIG-bench [1], MMLU-Pro [30], WikiText [22], and Tako [15].
@@
  • Uniform. Experts are uniformly distributed across all available devices. For instance, in our experimental setup with 4 GPUs, each device hosts 2 random experts per layer for the Mixtral model and 16 experts per layer for Deepseek-V2-Lite. This placement method aligns with the expert parallelism implementation in the Megatron-LM [28] project.
@@
  • SmartMoE [36]. While SmartMoE’s core focus is training optimization, our framework specifically implements SmartMoE’s placement module, which strategically distributes experts across GPUs based on real-time workload. The placement algorithm balances workload distribution to maintain uniform expert allocation across devices.
  • Expert Parallelism Load Balancer (EPLB) [3]. EPLB is a redundant expert strategy in DeepSeek-V3 [21] that duplicates high-load experts and heuristically distributes them to balance GPU workloads. Since the open-source implementation only supports homogeneous systems, we re-implement the algorithm within our framework to accommodate heterogeneous system configurations.
@@ IV-B System Performance
Inference Latency. As shown in Tab. II, robustness across datasets and server configurations highlights practical superiority of Prism. Uniform and SmartMoE achieve limited success due to inefficient resource utilization, while Redundance, though introducing expert duplication, suffers from higher latency due to suboptimal implementation. On the BigBench dataset, Prism reduces the average latency by 30.6% compared to EPLB, the second-best method for DeepSeek-V2-Lite. For Mixtral, it achieves a slight but consistent improvement over EPLB. On the MultiData dataset, Prism demonstrates a 10.9% reduction in latency compared to EPLB for DeepSeek-V2-Lite and a 2.0% improvement for Mixtral. Since EPLB is a load-balancing algorithm that does not account for cross-machine communication reduction, it underperforms compared to Prism. For the Mixtral model, where the number of experts is relatively small and server capacity is limited, the room for optimization is narrower, resulting in smaller performance gains for Prism.
Local Compute Ratio. To assess the impact of cross-machine communication on inference latency, we measured the local computation ratio as an indicator of communication overhead. Except for the Uniform and Redundance baselines, all methods adopt Prism’s migration strategy but differ in their expert placement strategies. As shown in Fig. 6, Prism consistently achieves higher local computation ratios. Notably, the advantage becomes more pronounced following explicit migration events triggered at 5 minutes (BigBench) and 10 minutes (MultiData).
@@ IV-C Large-scale Simulation
Simulation Setup. The simulator’s primary function orchestrates the inference process by: selecting the next layer for processing; updating server timestamps; recording completion times; computing inference latency. Through incremental time accumulation, it accurately models the temporal progression of the entire system. The simulator contains several key components: (1) Prompt Sequence Generator: We integrate a Poisson arrival time sequence with operational data collected from Prism, the dataset contains both expert selection patterns and token processing volumes. (2) Prompt Routing Generator: To simulate data flow under varying expert placement configurations. (3) Communication & Computation Time Consumption Estimator: We develop a linear model to predict processing time per token batch. (4) Time Stamp Calculator. This module captures both communication and computation events, timestamping each operation. (5) System Timeline Scheduler. The scheduler carefully interleaves communication and computation events according to our prescribed design.